Aviation Inc. is considering a new inventory system that will cost $375,00. The system is expected to generate $315,000 in year one, -$25,000 (negative) in year two, $110,000 in year three, and $150,000 in year four. Aviation's required rate of return is 10%. What is the MIRR (modified internal rate of return) of this project?
Answer #1

Aviation Inc. is considering a new inventory system that will cost $375,00. The system is expected to generate $315,000 in year one, -$25,000 (negative) in year two, $110,000 in year three, and $150,000 in year four. Aviation's required rate of return is 10%. What is the MIRR (modified internal rate of return) of this project?

MIRR = FV of inflows V PV of out flows

Year CF Discount or Compound Factor CF
0 $    -37,500.00 1/(1+0.1)^0= 1 1*-37500= $     -37,500.00
1 $ 3,15,000.00 (1+0.1)^1= 1.1 1.1*315000= $   3,46,500.00
2 $    -25,000.00 1/(1+0.1)^2= 0.826446281 0.826446280991735*-25000= $     -20,661.16
3 $      11,000.00 (1+0.1)^3= 1.331 1.331*11000= $       14,641.00
4 $ 1,50,000.00 (1+0.1)^4= 1.4641 1.4641*150000= $   2,19,615.00

Negative values are only an indicator of outflow, we use only the absolute values in the MIRR formula

MIRR = 346500 + 14641 + 219615 37500 + 20661

5, 80, 756.00 MIRR = 1 V 58,161.16

MIRR = 0.7776

Therefore MIRR = 77.76%